# Conditioning of Banach Space Valued Gaussian Random Variables: An Approximation Approach Based on Martingales

### v1

@@ 4 Examples with a Focus on the Gaussian Process Case
for $\mathrm{P}_{Y}$-almost all observed paths $g\in C(S)$ and all $s\in S$ and $t\in T$, where again the convergence is uniform in $s$, respectively $s$ and $t$.
@@ 5 Proofs of the Results of Section 3
Moreover, an $E$-valued, weakly measurable random variable $X$ is a Gaussian random variable, if and only if $\langle x^{\prime},X\rangle_{E^{\prime},E}$ is an $\mathbb{R}$-valued Gaussian random variable for all $x^{\prime}\in\mathfrak{D}$.
Proof of Theorem 5.1: Let us first consider the case of an $E$-valued and weakly measurable random variable $X:\Omega\to E$. We begin by assuming that $\mathrm{P}_{x^{\prime}}$ is a Gaussian measure for all $x^{\prime}\in\mathfrak{D}$. To this end, we fix an $x^{\prime}\in E^{\prime}$. Since the relative $w^{*}$-topology of $B_{E^{\prime}}$ is metrizable, see Theorem 2.1, there then exists a sequence $(x^{\prime}_{n})\subset\mathfrak{D}$ with $\langle x_{n}^{\prime},x\rangle\to\langle x^{\prime},x\rangle$ for all $x\in E$. Consequently, we have
@@
and this in turn shows that $\langle x_{n}^{\prime},X\rangle_{E^{\prime},E}\to\langle x^{\prime},X\rangle_{E^{\prime},E}$ in distribution. Now, all $\langle x^{\prime}_{n},X\rangle$ are $\mathbb{R}$-valued Gaussian random variables by assumption, and consequently Theorem C.4 ensures that $\langle x^{\prime},X\rangle$ is also a Gaussian random variable. This shows that $X$ is a Gaussian random variable. The converse implication is trivial.
Finally, the assertion for Gaussian measures follows from considering the random variable $X:=\operatorname{id}_{E}:E\to E$ on the probability space $(E,\sigma(E^{\prime}),\mathrm{P})$, since in this case we have $\mathrm{P}_{X}=\mathrm{P}$. ∎
@@
Let $E$ be a Banach space and $T:E^{\prime}\to E$ be a bounded, linear operator. Then we have
$\bigl{(}x^{\prime}\circ T\circ(x^{\prime})^{\prime}\bigr{)}(\langle t,\,\cdot\,\rangle_{\mathbb{R}})=\langle x^{\prime},Tx^{\prime}\rangle_{E^{\prime},E}\cdot t\,,\qquad\qquad x^{\prime}\in E^{\prime},t\in\mathbb{R},$
@@
Proof of Lemma 5.2: We fix an $x^{\prime}\in E^{\prime}$ and recall that the elements in $\mathbb{R}^{\prime}$ can be represented by $\mathbb{R}$ via the canonical isomorphism $t\mapsto\langle t,\,\cdot\,\rangle_{\mathbb{R}}$. Now observe that for $t\in\mathbb{R}$ and $x\in E$ we have
@@
Now, [23, Example 1.8] shows that ${\cal G}(X)$ is a Gaussian Hilbert space. However, this also directly follows from the observation that ${\cal G}(X)$ is a closed subspace of ${L_{2}(\mathrm{P})}$ by construction and from the fact that ${L_{2}(\mathrm{P})}$-limits of centered Gaussian random variables are centered Gaussian random variables, see also Theorem C.4.
Moreover, each $\langle x^{\prime},X\rangle_{E^{\prime},E}\in{\cal G}(X)$ is $\sigma(X)$-measurable, and therefore there is nothing to prove for the $\mathrm{P}$-equivalence classes of these functions. Let us now fix an arbitrary $f\in{\cal G}(X)$, so that our goal is to find a $\sigma(X)$-measurable $g:\Omega\to\mathbb{R}$ with $\mathrm{P}(\{f\neq g\})=0$. To this end, we first note that by the definition of ${\cal G}(X)$, there exists a sequence $(x^{\prime}_{n})\subset E^{\prime}$ with $f_{n}:=\langle x^{\prime}_{n},X\rangle_{E^{\prime},E}\to f$ in ${L_{2}(\mathrm{P})}$. Since there exists a $\mathrm{P}$-almost surely converging subsequence, see e.g. [27, Corollary 6.14], we may assume without loss of generality that we actually have the $\mathrm{P}$-almost sure convergence $f_{n}\to f$. Moreover, each $f_{n}$ is $\sigma(X)$-measurable, and therefore we have
@@
see e.g. [7, Lemma 2.1.7]. In addition, the $\mathrm{P}$-almost sure convergence $f_{n}\to f$ shows $\mathrm{P}(A)=1$. We define $g_{n}:=\boldsymbol{1}_{A}f_{n}$ and $g:\Omega\to\mathbb{R}$ by $g(\omega):=\lim f_{n}(\omega)$ for $\omega\in A$ and $g(\omega):=0$ for $\omega\in\Omega\setminus A$. Since each $g_{n}$ is $\sigma(X)$-measurable and $g_{n}(\omega)\to g(\omega)$ for all $\omega\in\Omega$ we then see that $g$ is $\sigma(X)$-measurable. Moreover, we have
$\mathrm{P}(\{f\neq g\})=\mathrm{P}(A\cap\{f\neq g\})+\mathrm{P}\bigl{(}(\Omega\setminus A)\cap\{f\neq g\})\bigr{)}=\mathrm{P}(A\cap\{f\neq\lim f_{n}\})=0\,,$
and hence we have found the desired $\sigma(X)$-measurable representative of the $\mathrm{P}$-equivalence class of $f$.
@@
ii). The inclusion $\supset$” follows from i) by considering the projections $\pi_{E}:E\times F\to E$ and $\pi_{F}:E\times F\to F$ and the fact that ${\cal G}(X,Y)$ is a closed linear subspace of ${L_{2}(\mathrm{P})}$. To prove the converse implication, we fix an $z^{\prime}\in(E\times F)^{\prime}$. As discussed in front of Definition 2.4, there then exist $x^{\prime}\in E^{\prime}$ and $y^{\prime}\in F^{\prime}$ such that $z^{\prime}(x,y)=x^{\prime}(x)+y^{\prime}(y)$ for all $x\in E$ and $y\in F$, and this in turn implies
@@
Our next goal is to describe the relationship between $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,)$ and the just considered $\mathrm{P}_{X_{i}|Y}(\,\cdot\,|\,\cdot\,)$ with the help of Theorem D.2, which describes regular conditional probabilities under transformations. To this end, we set $T_{0}:=E$ and $T_{1}:=\mathbb{R}$, as well as $U_{0}:=U_{1}:=\mathbb{R}^{n}$ and ${\cal A}_{0}:={\cal A}_{1}:=\mathfrak{B}^{n}$. Moreover, we consider the maps $\Phi:=x_{i}^{\prime}$ and $\Psi:=\operatorname{id}_{\mathbb{R}^{n}}$ as well as the measures $\mathrm{P}^{(0)}:=\mathrm{P}_{(X,Y)}$ and $\mathrm{P}^{(1)}:=\mathrm{P}_{(X_{i},Y)}=\mathrm{P}^{(1)}_{(\Phi,\Psi)}$. By Theorem D.2 there then exists an $N_{i}^{\prime\prime}\in\mathfrak{B}^{n}$ with $\mathrm{P}_{Y}(N_{i}^{\prime\prime})=0$ such that for all $y\in\mathbb{R}^{n}\setminus N_{i}^{\prime\prime}$ we have
@@
Obviously, this gives $N\in\mathfrak{B}^{n}$ with $\mathrm{P}_{Y}(N)=0$, and by construction (36) holds for all $y\in\mathbb{R}^{n}\setminus N$ and $x^{\prime}_{i}\in\mathfrak{D}$. Let us fix a $y\in\mathbb{R}^{n}\setminus N$. Then we have just seen in (36) that for all $x_{i}^{\prime}\in\mathfrak{D}$ the image measure $(\mathrm{P}_{X|Y}(\,\cdot\,|y))_{x^{\prime}_{i}}$ of $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ under $x_{i}^{\prime}$ is Gaussian, and by Theorem 5.1 we can thus conclude that $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ is Gaussian.
@@
where we have also used a projection formula for covariances (B.6), the basic transformation formula for $\mathbb{R}$-valued integrals, and (37). Combining this with (35) we thus find
@@
for all $x_{i}^{\prime}\in\mathfrak{D}$ and $y\in\mathbb{R}^{n}\setminus N$. Now the left hand side of this identity is $w^{*}$-continuous on $B_{E^{\prime}}$ by Theorem B.3, and the right-hand side is $w^{*}$-continuous on $B_{E^{\prime}}$ by Theorem B.3 and Lemma B.4. Since $\mathfrak{D}$ is $w^{*}$-dense in $B_{E^{\prime}}$, we thus find
@@
for all $x^{\prime}\in E^{\prime}$. Moreover, $(\operatorname{cov}Y)^{\dagger}:\mathbb{R}^{n}\to(\mathbb{R}^{n})^{\prime}$ is symmetric and non-negative, see the discussion around (B.8), and therefore Lemma B.4 shows that $R_{X,Y}$ is an abstract covariance operator. This in turn shows that
@@
$\mu_{X|Y}(Y(\omega))=\int_{E}x\,\mathrm{P}_{X|Y}(\,\mathrm{d}x|Y(\omega))$ $=\int_{E}\boldsymbol{1}_{\mathbb{R}^{n}\setminus N}(Y(\omega))\cdot\bigl{(}(g\circ\pi_{E})(x,Y(\omega))\bigr{)}\,\mathrm{d}\mathrm{P}_{X|Y}(\,\mathrm{d}x|Y(\omega))$
@@
$\int_{\Omega}\boldsymbol{1}_{A_{0}}\cdot X\,\mathrm{d}\mathrm{P}=\int_{\Omega}\boldsymbol{1}_{A}\circ Y\cdot X\,\mathrm{d}\mathrm{P}$ $=\int_{\Omega}\boldsymbol{1}_{A}\circ\pi_{\mathbb{R}^{n}}(X,Y)\cdot g\circ\pi_{E}(X,Y)\,\mathrm{d}\mathrm{P}$
$=\int_{E\times\mathbb{R}^{n}}\boldsymbol{1}_{E\times A}\cdot g\circ\pi_{E}\,\mathrm{d}\mathrm{P}_{(X,Y)}\,.$
@@
$Ty:=\mathbb{E}X+\operatorname{cov}(X,Y)(\operatorname{cov}Y)^{\dagger}(y-\mathbb{E}Y)\,,\qquad\qquad y\in\mathbb{R}^{n},$
@@
by the definition of $T$ and the already established formula for $\mu_{X|Y}(y)$. Since $E=\mathbb{R}$ we further have $K_{X|Y}=\operatorname{var}(\mathrm{P}_{X|Y}(\,\cdot\,|y))$ for all $y\in\mathbb{R}^{n}\setminus N$ and combining this with (41) and (40) leads to
@@
Finally, note that our definition of $Z$ gives $Z-\mathbb{E}Z=Z-\mathbb{E}X=\operatorname{cov}(X,Y)(\operatorname{cov}Y)^{\dagger}(Y-\mathbb{E}Y)$. Moreover, $Y-\mathbb{E}Y$ is a centered Gaussian random variable and $\operatorname{cov}(X,Y)(\operatorname{cov}Y)^{\dagger}:\mathbb{R}^{n}\to E$ is bounded and linear. For $x^{\prime}\in E^{\prime}$ we then have $y^{\prime}:=x^{\prime}\circ\operatorname{cov}(X,Y)\circ(\operatorname{cov}Y)^{\dagger}\in(\mathbb{R}^{n})^{\prime}$, and this gives
$\langle x^{\prime},Z-\mathbb{E}Z\rangle_{E^{\prime},E}=\bigl{(}x^{\prime}\circ\operatorname{cov}(X,Y)\circ(\operatorname{cov}Y)^{\dagger}\bigr{)}(Y-\mathbb{E}Y)=\langle y^{\prime},Y-\mathbb{E}Y\rangle_{(\mathbb{R}^{n})^{\prime},\mathbb{R}^{n}}\in{\cal G}(Y)\,.$
Taking the ${L_{2}(\mathrm{P})}$-closure then yields the assertion. ∎
@@
In this subsection we elaborate on filtering sequences and prove Proposition 3.3. We begin with the following lemma that considers the behavior of generated $\sigma$-algebras under some natural operations in $F^{\prime}$.
Lemma 5.5.
Let $F$ be a Banach space and $B\subset F^{\prime}$ be non-empty. Then we have
Proof of Lemma 5.5: We first prove the equation $\sigma(B)=\sigma(\operatorname{span}B)$. Here, the inclusion $\subset$” is obvious. To establish the converse, we fix an $y^{\prime}\in\operatorname{span}B$. Then there exist an $n\geq 1$ as well as $\alpha_{1},\dots,\alpha_{n}\in\mathbb{R}$, and $y_{1}^{\prime},\dots,y_{n}^{\prime}\in B$ with $y^{\prime}=\alpha_{1}y_{1}^{\prime}+\dots+\alpha_{n}y_{n}^{\prime}$. Clearly, each $y_{i}^{\prime}:F\to\mathbb{R}$ is $(\sigma(B),\mathfrak{B})$-measurable, and hence $y^{\prime}$ is also $(\sigma(B),\mathfrak{B})$-measurable. This gives $\sigma(y^{\prime})\subset\sigma(B)$, and since $y^{\prime}\in\operatorname{span}B$ was chosen arbitrarily, we find $\sigma(\operatorname{span}B)\subset\sigma(B)$.
The third equation is a direct consequence of the first two equations. ∎
Lemma 5.6.
  1. i)
  2. ii)
Proof of Lemma 5.6: i). We first note that in view of Lemma 5.5 we only need to show $\sigma(A)=\sigma(\mathfrak{C}_{A})$. To establish this identity, let $a_{1}^{\prime},\dots,a_{n}^{\prime}:F\to\mathbb{R}$ be the coordinate functions of $A$, that is $\mathfrak{C}_{A}=\{a_{1}^{\prime},\dots,a_{n}^{\prime}\}$. Then, [27, Theorem 1.90] shows that for a given $\sigma$-algebra ${\cal A}^{\prime}$ on $F$, the map $A:F\to\mathbb{R}^{n}$ is $({\cal A}^{\prime},\mathfrak{B}^{n})$-measurable if and only if all $a_{1}^{\prime},\dots,a_{n}^{\prime}$ are $({\cal A}^{\prime},\mathfrak{B})$-measurable.
ii). Let $e_{1},\dots,e_{n}$ be the standard ONB of $\mathbb{R}^{n}$. Then $A$ is given by $A=\sum_{i=1}^{n}a_{i}^{\prime}\otimes e_{i}$, and therefore (B.14) shows that its adjoint $A^{\prime}$ is given by
where the first identity can be found in e.g. [32, Lemma 3.1.16] or [11, Proposition VI.1.8], and the second identity can be found in e.g. [32, Proposition 2.6.6]. Moreover, since the subspace $\operatorname{ran}A^{\prime}$ of $F^{\prime}$ is finite dimensional, it is $w^{*}$-closed, see e.g. [32, Corollary 2.2.32] or [37, Theorem 1.21]. Combining the latter with (42) yields $(\ker A)^{\perp\perp}=\operatorname{ran}A^{\prime}$. ∎
Proposition 5.7.
$F_{\infty}^{\prime}:=\bigcup_{n\geq 1}\operatorname{span}\mathfrak{C}_{A_{n}}\,.$
$=\sigma\biggl{(}\bigcup_{n\geq 1}\sigma(A_{n})\biggr{)}\,,$
Applying Proposition 5.7 then yields the assertion. ∎
@@
where in the second and the last step we used an elementary formula for computing the $\sigma$-algebra generated by a single map, see e.g. [27, Theorem 1.78]. Since $(\sigma(A_{n}))$ is a filtration, so is $(\sigma(Y_{n}))$. To establish the formula for $\sigma(Y)$, we write ${\cal E}:=\bigcup_{n\geq 1}\sigma(A_{n})$. By assumption we know that $\sigma({\cal E})=\sigma(F^{\prime})$, and hence we obtain
@@
where in the third step we used another well-known fact on generated $\sigma$-algebras, see e.g. [27, Theorem 1.81] and its proof. ∎
@@
Let $E$ be a separable Banach space. Then a set ${\cal M}$ of probability measures on $(E,\mathfrak{B}(E))$ is called uniformly tight, if for all $\varepsilon>0$ there exists a compact $K\subset E$ with $\mu(E\setminus K)\leq\varepsilon$ for all $\mu\in{\cal M}$. Recall that given a probability measure $\mu$ on $(E,\mathfrak{B}(E))$, Ulam’s theorem, see e.g. [4, Lemma 26.2] or [46, Theorem 3.1 in Chapter I.3.2], shows that $\mu$ a Radon measure in the sense of inner regularity, that is
@@
for all $B\in\mathfrak{B}(E)$. In particular, $\{\mu\}$ is (uniformly) tight. Here we emphasize that we use the term “Radon measure” in the sense of [46, p. 28] and [6, p. 97], as we will need this definition also in the proof of Theorem C.6 when applying results from [46] and [6]. Consequently, every finite set ${\cal M}$ of probability measures on $(E,\mathfrak{B}(E))$ is uniformly tight and more generally, the finite union of uniformly tight sets is also uniformly tight.
The following lemma, which provides a simple characterization of uniform tightness, is taken from [8, Example 8.6.5]. In view of its formulation we note that [8, Example 8.6.5] additionally lists the condition $\mu(\{V=\infty\})=0$ for all $\mu\in{\cal M}$. However, this condition is implied by the finiteness of the integrals in ii), and hence we omitted it.
@@
In addition, for all Banach spaces $G$ and all $\mathrm{P}_{X}$-Bochner integrable $g:E\to G$ there exists an $N_{g}\in\sigma(F^{\prime})$ with $\mathrm{P}_{Y}(N_{g})=0$ such that for all $y\in F\setminus N_{g}$ we have $g\in{{\cal L}_{1}(\mathrm{P}_{X|Y}(\,\cdot\,|y),G)}$ and
$g\in{{\cal L}_{1}(\mathrm{P}_{X|Y_{n}}(\,\cdot\,|A_{n}(y)),G)}\,,\qquad\qquad n\geq 1,$ (45)

### v3

@@ 5 Conditioning Gaussian Processes with Continuous Paths
for $\mathrm{P}_{Y}$-almost all observed paths $g\in C(S)$ and all $s\in S$ and $t\in T$, where again the convergence is uniform in $s$, respectively $s$ and $t$. Note that these findings substantially generalize the state-of-the-art established in [30] in the sense that neither $S=T$ nor additional assumptions on the observational model are needed.
@@ 6 Proofs of the Results of Section 3
Moreover, an $E$-valued random variable $X$ is a Gaussian random variable, if and only if $\langle x^{\prime},X\rangle_{E^{\prime},E}$ is an $\mathbb{R}$-valued Gaussian random variable for all $x^{\prime}\in\mathfrak{D}$.
Proof of Theorem 6.1: Let us first consider the case of an $E$-valued random variable $X:\Omega\to E$. We begin by assuming that $\langle x^{\prime},X\rangle_{E^{\prime},E}$ is a Gaussian random variable for all $x^{\prime}\in\mathfrak{D}$. Let us now fix an $x^{\prime}\in E^{\prime}$. Since the relative $w^{*}$-topology of $B_{E^{\prime}}$ is metrizable, see Theorem 2.1, there then exists a sequence $(x^{\prime}_{n})\subset\mathfrak{D}$ with $\langle x_{n}^{\prime},x\rangle\to\langle x^{\prime},x\rangle$ for all $x\in E$. Consequently, we have
@@
and this in turn shows that $\langle x_{n}^{\prime},X\rangle_{E^{\prime},E}\to\langle x^{\prime},X\rangle_{E^{\prime},E}$ in distribution. Now, all $\langle x^{\prime}_{n},X\rangle$ are $\mathbb{R}$-valued Gaussian random variables by assumption, and consequently Theorem C.4 ensures that $\langle x^{\prime},X\rangle$ is also a Gaussian random variable. Consequently, $X$ is a Gaussian random variable. The converse implication is trivial.
Finally, the assertion for Gaussian measures follows from considering the random variable $X:=\operatorname{id}_{E}:E\to E$ on the probability space $(E,\sigma(E^{\prime}),\mathrm{P})$, since in this case we have $\mathrm{P}_{X}=\mathrm{P}$ and $\mathrm{P}_{\langle x^{\prime},X\rangle_{E^{\prime},E}}=\mathrm{P}_{x^{\prime}}$. ∎
@@
Let $E$ be a Banach space, $T:E^{\prime}\to E$ be a bounded, linear operator, and $x^{\prime}\in E^{\prime}$. Then we have
$\bigl{(}x^{\prime}\circ T\circ(x^{\prime})^{\prime}\bigr{)}(\langle t,\,\cdot\,\rangle_{\mathbb{R}})=\langle x^{\prime},Tx^{\prime}\rangle_{E^{\prime},E}\cdot t\,,\qquad\qquad t\in\mathbb{R},$
@@
Proof of Lemma 6.2: Recall that the elements in $\mathbb{R}^{\prime}$ can be represented by $\mathbb{R}$ via the canonical isomorphism $t\mapsto\langle t,\,\cdot\,\rangle_{\mathbb{R}}$. Now observe that for $t\in\mathbb{R}$ and $x\in E$ we have
@@
Now, [26, Example 1.8] shows that ${\cal G}(X)$ is a Gaussian Hilbert space. However, this also directly follows from the observation that ${\cal G}(X)$ is a closed subspace of ${L_{2}(\mathrm{P})}$ by construction and from the fact that ${L_{2}(\mathrm{P})}$-limits of centered Gaussian random variables are centered Gaussian random variables, see also Theorem C.4.
Let us now find a $\sigma(X)$-measurable representative. Clearly, each $\langle x^{\prime},X\rangle_{E^{\prime},E}\in{\cal G}(X)$ is $\sigma(X)$-measurable, and therefore there is nothing to prove for the $\mathrm{P}$-equivalence classes of these functions. Let us now fix an arbitrary $f\in{\cal G}(X)$, so that our goal is to find a $\sigma(X)$-measurable $g:\Omega\to\mathbb{R}$ with $\mathrm{P}(\{f\neq g\})=0$. To this end, we first note that by the definition of ${\cal G}(X)$, there exists a sequence $(x^{\prime}_{n})\subset E^{\prime}$ with $f_{n}:=\langle x^{\prime}_{n},X\rangle_{E^{\prime},E}\to f$ in ${L_{2}(\mathrm{P})}$. Since there exists a $\mathrm{P}$-almost surely converging subsequence, we may assume without loss of generality that we actually have the $\mathrm{P}$-almost sure convergence $f_{n}\to f$. Moreover, each $f_{n}$ is $\sigma(X)$-measurable, and therefore we have
@@
see e.g. [9, Lemma 2.1.7]. In addition, the $\mathrm{P}$-almost sure convergence $f_{n}\to f$ shows $\mathrm{P}(A)=1$. We define $g_{n}:=\boldsymbol{1}_{A}f_{n}$ and $g:\Omega\to\mathbb{R}$ by $g(\omega):=\lim_{n\to\infty}f_{n}(\omega)$ for $\omega\in A$ and $g(\omega):=0$ for $\omega\in\Omega\setminus A$. Since each $g_{n}$ is $\sigma(X)$-measurable and $g_{n}(\omega)\to g(\omega)$ for all $\omega\in\Omega$ we then see that $g$ is $\sigma(X)$-measurable. Moreover, we have
$\mathrm{P}(\{f\neq g\})=\mathrm{P}(A\cap\{f\neq g\})+\mathrm{P}\bigl{(}(\Omega\setminus A)\cap\{f\neq g\})\bigr{)}=\mathrm{P}(A\cap\{f\neq\lim_{n\to\infty}f_{n}\})=0\,,$
and hence we have found the desired $\sigma(X)$-measurable representative $g$ of the $\mathrm{P}$-equivalence class of $f$.
@@
ii). For the inclusion $\supset$” we consider the coordinate projection $\pi_{E}:E\times F\to E$. Then we have $\pi_{E}(X,Y)=X$ and therefore i) gives ${\cal G}(X)={\cal G}(\pi_{E}(X,Y))\subset{\cal G}(X,Y)$. Since we analogously find ${\cal G}(Y)\subset{\cal G}(X,Y)$, we conclude that ${\cal G}(X)+{\cal G}(Y)\subset{\cal G}(X,Y)$. Taking the ${L_{2}(\mathrm{P})}$-closure then yields the desired inclusion since ${\cal G}(X,Y)$ is a closed linear subspace of ${L_{2}(\mathrm{P})}$. To prove the converse implication, we fix an $z^{\prime}\in(E\times F)^{\prime}$. As discussed in front of Definition 2.4, there then exist $x^{\prime}\in E^{\prime}$ and $y^{\prime}\in F^{\prime}$ such that $z^{\prime}(x,y)=x^{\prime}(x)+y^{\prime}(y)$ for all $x\in E$ and $y\in F$, and this in turn implies
@@
Our next goal is to describe the relationship between $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,)$ and the just considered $\mathrm{P}_{X_{i}|Y}(\,\cdot\,|\,\cdot\,)$ with the help of Theorem D.2, which describes regular conditional probabilities under transformations. To this end, we set $T_{0}:=E$ and $T_{1}:=\mathbb{R}$, as well as $U_{0}:=U_{1}:=\mathbb{R}^{n}$ and ${\cal A}_{0}:={\cal A}_{1}:=\mathfrak{B}^{n}$. Moreover, we consider the maps $\Phi:=x_{i}^{\prime}$ and $\Psi:=\operatorname{id}_{\mathbb{R}^{n}}$ as well as the measures $\mathrm{P}^{(0)}:=\mathrm{P}_{(X,Y)}$ and $\mathrm{P}^{(1)}:=\mathrm{P}_{(X_{i},Y)}=\mathrm{P}^{(1)}_{(\Phi,\Psi)}$. By Theorem D.2 there then exists an $N_{i}^{\prime\prime}\in\mathfrak{B}^{n}$ with $\mathrm{P}_{Y}(N_{i}^{\prime\prime})=\mathrm{P}_{\mathbb{R}^{n}}^{(0)}(N_{i}^{\prime\prime})=0$ such that for all $y\in\mathbb{R}^{n}\setminus N_{i}^{\prime\prime}$ we have
@@
Obviously, this gives $N\in\mathfrak{B}^{n}$ with $\mathrm{P}_{Y}(N)=0$, and by construction (39) holds for all $y\in\mathbb{R}^{n}\setminus N$ and $x^{\prime}_{i}\in\mathfrak{D}$.
Let us now fix a $y\in\mathbb{R}^{n}\setminus N$. Then we have just seen in (39) that for all $x_{i}^{\prime}\in\mathfrak{D}$ the image measure $(\mathrm{P}_{X|Y}(\,\cdot\,|y))_{x^{\prime}_{i}}$ of $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ under $x_{i}^{\prime}$ is Gaussian, and by Theorem 6.1 we can thus conclude that $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ is Gaussian.
@@
where we have also used a projection formula for covariances (B.6), the basic transformation formula for $\mathbb{R}$-valued integrals, (40), and (B.1). Combining this with (38) we thus find
@@
for all $x_{i}^{\prime}\in\mathfrak{D}$ and $y\in\mathbb{R}^{n}\setminus N$. Now the left hand side of this identity is $w^{*}$-continuous on $B_{E^{\prime}}$ by Theorem B.3. Moreover, $(\operatorname{cov}Y)^{\dagger}:\mathbb{R}^{n}\to(\mathbb{R}^{n})^{\prime}$ is symmetric and non-negative, see the discussion around (B.8). Consequently, the right-hand side is also $w^{*}$-continuous on $B_{E^{\prime}}$ by Theorem B.3 and Lemma B.4. Since $\mathfrak{D}$ is $w^{*}$-dense in $B_{E^{\prime}}$, we thus find
@@
for all $x^{\prime}\in E^{\prime}$. Moreover, since $(\operatorname{cov}Y)^{\dagger}:\mathbb{R}^{n}\to(\mathbb{R}^{n})^{\prime}$ is symmetric and non-negative, Lemma B.4 shows that $R_{X,Y}$ is an abstract covariance operator, and this implies that
@@
$Z(\omega)=\mu_{X|Y}(Y(\omega))=\int_{E}x\,\mathrm{P}_{X|Y}(\,\mathrm{d}x|Y(\omega))$ $=\int_{E}\boldsymbol{1}_{\mathbb{R}^{n}\setminus N}(Y(\omega))\cdot\bigl{(}(g\circ\pi_{E})(x,Y(\omega))\bigr{)}\,\mathrm{d}\mathrm{P}_{X|Y}(\,\mathrm{d}x|Y(\omega))$
@@
$\int_{\Omega}\boldsymbol{1}_{A_{0}}\cdot X\,\mathrm{d}\mathrm{P}=\int_{\Omega}(\boldsymbol{1}_{A}\circ Y)\cdot X\,\mathrm{d}\mathrm{P}$ $=\int_{\Omega}\bigl{(}\boldsymbol{1}_{A}\circ\pi_{\mathbb{R}^{n}}(X,Y)\bigr{)}\cdot\bigl{(}g\circ\pi_{E}(X,Y)\bigr{)}\,\mathrm{d}\mathrm{P}$
$=\int_{E\times\mathbb{R}^{n}}\boldsymbol{1}_{E\times A}\cdot(g\circ\pi_{E})\,\mathrm{d}\mathrm{P}_{(X,Y)}\,.$
@@
$Ty:=\mathbb{E}X+\operatorname{cov}(X,Y)(\operatorname{cov}Y)^{\dagger}(y-\mathbb{E}Y)\,,\qquad\qquad y\in\mathbb{R}^{n},$ (43)
@@
by the definition (43) of $T$ and the already established formula for $\mu_{X|Y}(y)$. Since $E=\mathbb{R}$ we further have $K_{X|Y}=\operatorname{var}(\mathrm{P}_{X|Y}(\,\cdot\,|y))$ for all $y\in\mathbb{R}^{n}\setminus N$ and combining this with (45) and (44) leads to
@@
Finally, for the proof of ${\cal G}(Z)\subset{\cal G}(Y)$, we consider the bounded linear operator $A:=\operatorname{cov}(X,Y)(\operatorname{cov}Y)^{\dagger}:\mathbb{R}^{n}\to E$. The definition of $Z$ then gives
$Z-\mathbb{E}Z=Z-\mathbb{E}X=\operatorname{cov}(X,Y)(\operatorname{cov}Y)^{\dagger}(Y-\mathbb{E}Y)=A(Y-\mathbb{E}Y)\,.$
By Lemma 6.4 we thus conclude ${\cal G}(Z)={\cal G}(Z-\mathbb{E}Z)\subset{\cal G}(Y-\mathbb{E}Y)={\cal G}(Y)$. ∎
@@
where in the second and the last step we used an elementary formula for computing the $\sigma$-algebra generated by a single map, see e.g. [29, Theorem 1.78]. Since $(\sigma(A_{n}))$ is a filtration, so is $(\sigma(Y_{n}))$.
To establish the formula for $\sigma(Y)$, we write ${\cal E}:=\bigcup_{n\geq 1}\sigma(A_{n})$. By assumption we know that $\sigma({\cal E})=\sigma(F^{\prime})$, and hence we obtain
@@
where in the third step we used another well-known fact on generated $\sigma$-algebras, see e.g. [29, Theorem 1.81] and its proof. ∎
@@
Let $E$ be a separable Banach space. Then a set ${\cal M}$ of probability measures on $(E,\mathfrak{B}(E))$ is called uniformly tight, if for all $\varepsilon>0$ there exists a compact $K\subset E$ with $\mu(E\setminus K)\leq\varepsilon$ for all $\mu\in{\cal M}$. To quickly discuss the definition recall that given a probability measure $\mu$ on $(E,\mathfrak{B}(E))$, Ulam’s theorem, see e.g. [6, Lemma 26.2] or [50, Theorem 3.1 in Chapter I.3.2], shows that $\mu$ is a Radon measure in the sense of inner regularity, that is
@@
for all $B\in\mathfrak{B}(E)$. In particular, $\{\mu\}$ is (uniformly) tight. Here we emphasize that we use the term “Radon measure” in the sense of [50, p. 28], [8, p. 97] and [10, Definition 7.1.1], as we will need this definition also in the proof of Theorem C.6 when applying results from [50] and [8]. By (46) we conclude that every finite set ${\cal M}$ of probability measures on $(E,\mathfrak{B}(E))$ is uniformly tight. Similarly, the finite union of uniformly tight sets is also uniformly tight.
The following lemma, which provides a simple characterization of uniform tightness, is taken from [10, Example 8.6.5]. In view of its formulation we note that [10, Example 8.6.5] additionally lists the condition $\mu(\{V=\infty\})=0$ for all $\mu\in{\cal M}$. However, this condition is implied by the finiteness of the integrals in ii), and hence we omitted it.
@@
With these preparations we can now present the main result of this section, which, roughly speaking, translates martingale convergence to convergence of integrals taken with respect to regular conditional probabilities.
@@
In addition, for all Banach spaces $G$ and all $\mathrm{P}_{X}$-Bochner integrable $g:E\to G$ there exists an $N_{g}\in\sigma(F^{\prime})$ with $\mathrm{P}_{Y}(N_{g})=0$ such that for all $y\in F\setminus N_{g}$ we have
$g\in{{\cal L}_{1}(\mathrm{P}_{X|Y}(\,\cdot\,|y),G)}\qquad\qquad\mbox{ and }\qquad\qquad g\in{{\cal L}_{1}(\mathrm{P}_{X|Y_{n}}(\,\cdot\,|A_{n}(y)),G)}\,,\qquad\qquad n\geq 1,$ (47)